# DART: A Diffusion-Based Autoregressive Motion Model for Real-Time Text-Driven Motion Control

### v1

@@ Abstract
Text-conditioned human motion generation, which allows for user interaction through natural language, has become increasingly popular. Existing methods typically generate short, isolated motions based on a single input sentence. However, human motions are continuous and can extend over long periods, carrying rich semantics. Creating long, complex motions that precisely respond to streams of text descriptions, particularly in an online and real-time setting, remains a significant challenge. Furthermore, incorporating spatial constraints into text-conditioned motion generation presents additional challenges, as it requires aligning the motion semantics specified by text descriptions with geometric information, such as goal locations and 3D scene geometry. To address these limitations, we propose DART, a Diffusion-based Autoregressive motion primitive model for Real-time Text-driven motion control. Our model, DART, effectively learns a compact motion primitive space jointly conditioned on motion history and text inputs using latent diffusion models. By autoregressively generating motion primitives based on the preceding history and current text input, DART enables real-time, sequential motion generation driven by natural language descriptions. Additionally, the learned motion primitive space allows for precise spatial motion control, which we formulate either as a latent noise optimization problem or as a Markov decision process addressed through reinforcement learning. We present effective algorithms for both approaches, demonstrating our model’s versatility and superior performance in various motion synthesis tasks. Experiments show our method outperforms existing baselines in motion realism, efficiency, and controllability.
@@ 1 Introduction
Text-conditioned human motion generation has gained increasing popularity in recent years for flexible user interaction via natural languages. Existing text-conditioned motion models  (Tevet et al., 2023; Guo et al., 2023; Zhang et al., 2023a; Guo et al., 2022; Jiang et al., 2024a) primarily focus on generating standalone short motions from a single descriptive sentence, thereby overlooking the inherently sequential nature of human motion. These methods fail to generate long and complex motions consisting of multiple segments with distinct conditions of action descriptions. FlowMDM (Barquero et al., 2024) is the state-of-the-art temporal motion composition method, capable of generating complex, continuous motions while providing control over the desired actions and their durations. However, FlowMDM is an offline method that requires prior knowledge of the entire action timeline and has a slow generation speed, making it unsuitable for online and real-time applications.
@@
In addition to text-based semantic control, generating human motion within spatial constraints and achieving specific goals—such as reaching a keyframe body pose, following the joint trajectory, or interacting with objects-offers broad applications but introduces additional complex challenges. Recent works (Shafir et al., 2024; Karunratanakul et al., 2024; Xie et al., 2024) have sought to integrate text-conditioned motion models with spatial control capabilities. However, they still face challenges in simultaneously harmonizing spatial control, motion quality, and semantic alignment with text. Moreover, these approaches are typically restricted to controlling isolated short motions in an offline setting. Conversely, interactive character control (Kovar et al., 2008; Holden et al., 2015; Ling et al., 2020; Peng et al., 2022) has been a longstanding focus in computer graphics, with a primary emphasis on achieving motion realism and responsiveness to interactive control signals. However, most of these works lack support for text-conditioned semantic control and are limited by being trained on small, curated datasets. Incorporating text-conditioned motion generation could provide a novel, efficient language interface for animators and everyday users to control the characters, reducing the effort required when specifying spatial control signals is challenging or tedious.
@@
To address these limitations, we propose DART, a diffusion-based autoregressive motion primitive model for real-time text-conditioned motion composition. Moreover, the compact and expressive motion space of DART provides a foundation for integrating precise spatial control through latent space optimization or reinforcement learning-based algorithms. DART features three key components.
@@
First, DART represents human motion as a collection of motion primitives (Zhang & Tang, 2022), which are autoregressive representations consisting of overlapping short motion segments tailored for online generation and control. These short primitives also provide a clearer alignment with atomic action semantics compared to longer sequences, enabling effective learning of a text-conditioned motion space. By focusing on shorter primitives, DART avoids the complexities and extensive data demands of modeling entire motion sequences, allowing for high-quality motion generation with only a few diffusion steps.
@@
Next, DART learns a text-conditioned motion primitive space from large-scale data using a latent diffusion architecture, which contains a denoiser network for generating motions in a compact latent motion space, and a decoder network for reconstructing the detailed motions. Leveraging this learned latent motion space, DART employs an autoregressive rollout to synthesize motion sequences from real-time text inputs, enabling the efficient generation of motions of arbitrary length. Compared with the offline temporal motion composition method FlowMDM (Barquero et al., 2024), DART provides real-time response and ten times faster generation speed.
@@
Lastly, the versatile and powerful motion space learned in DART enables precise spatial control through latent space optimization or RL-based policy learning. We present effective learning and optimization algorithms that explore the learned DART motion space to obtain motion sequences that precisely follow both textual and spatial constraints. We evaluate DART across various motion synthesis tasks, including generating long, continuous sequences from sequential text prompts, in-between motion generation, scene-conditioned motion, and goal-reaching synthesis. The experimental results show that DART is a simple, unified and highly effective motion model, consistently outperforming or matching the performance of baselines.
@@ 2 Related Works
Conditional Motion Generation. Generating realistic and diverse human motions is a long-standing challenge in computer vision and graphics. Apart from generating highly realistic human motions (Kovar et al., 2008; Holden et al., 2020; Büttner & Clavet, 2015; Zinno, 2019), conditional generation is another important factor that aligns motion generation with human intentions and various application constraints. Text-conditioned motion generation (Tevet et al., 2023; Zhang et al., 2022; Petrovich et al., 2022; Guo et al., 2022; Jiang et al., 2024a; Zhang et al., 2023a) has become increasingly popular since it allows users to modulate motion generation with flexible natural languages. Audio and speech-driven motion synthesis methods (Alexanderson et al., 2023; Tseng et al., 2022; Siyao et al., 2022; Ao et al., 2022; 2023) have also made significant progress recently. Moreover, there exist many applications that require spatial awareness and precise control in motion generation, such as interactive character control (Ling et al., 2020; Peng et al., 2022; Starke et al., 2022; Luo et al., 2023; Starke et al., 2024), human-scene interactions(Hassan et al., 2021; Starke et al., 2019; Zhao et al., 2022; 2023; Li et al., 2024; Xu et al., 2023; Jiang et al., 2024b; Wang et al., 2024; Liu et al., 2023; Li et al., 2023a), and human-human(noid) interactions (Liang et al., 2024; Zhang et al., 2023c; Christen et al., 2023; Cheng et al., 2024; Shan et al., 2024). Synthesizing high-quality motions with precise spatial control remains challenging, and DART is a step toward a general and efficient motion model that supports precise control tasks.
@@
Diffusion Generative Models. Denoising Diffusion Models (Ho et al., 2020; Song et al., 2020a; b) are generative models that learn to predict clean data samples by gradually annealing the noise from a standard Gaussian sample. Diffusion models have achieved unprecedented performances in many generation tasks including images, videos, and 3D human motions (Tevet et al., 2023; Rombach et al., 2022; Ho et al., 2022). Diffusion models can accept flexible conditions to modulate the generation, such as text prompts, images, audio, and 3D objects (Rombach et al., 2022; Tevet et al., 2023; Alexanderson et al., 2023; Tseng et al., 2022; Zhang et al., 2023b; Xu et al., 2023; Li et al., 2023b). Most existing diffusion-based motion generation methods focus on offline generations (Tevet et al., 2023; Barquero et al., 2024; Chen et al., 2023; Karunratanakul et al., 2024; Cohan et al., 2024; Dai et al., 2024) producing short, isolated motion sequences at once while neglecting the autoregressive nature of human motions. Among these methods, DNO (Karunratanakul et al., 2024) is closely related to our optimization-based control approach, as both use diffusion noises as the latent space for editing and control. However, the key distinction lies in our latent motion primitive-based diffusion model, which, unlike DNO’s diffusion model trained with full motion sequences in raw space, achieves superior performance in harmonizing spatial control with text semantic alignment during experiments. There are also works that incorporate history conditions in diffusion-based motion prediction. Recent works (Xu et al., 2023; Jiang et al., 2024b) use history-conditioned diffusion models to autoregressive generate human-scene interaction sequences. Van Wouwe et al. (2024) uses autoregressive diffusion for real-time human motion reconstruction from IMU sensors. Shi et al. (2024) and Chen et al. (2024) adapt diffusion models for real-time character motion generation and control. While Shi et al. (2024) learns character control policies using diffusion noises as the action space, akin to our reinforcement-learning-based control, their method focuses on single-frame autoregressive generation and lacks support for text conditions, which offer a compact and powerful interface for users to control character behaviors. In contrast, DART is an efficient and general motion model that scales effectively to large motion datasets with text annotations (Mahmood et al., 2019; Punnakkal et al., 2021). DART supports natural language interfacing and provides a versatile foundation for various motion generation tasks with spatial control.
@@ 3.1 Preliminaries
Autoregressive Motion Primitive Representation. We model long-term human motions as the sequential composition of motion primitives (Zhang & Tang, 2022) with overlaps for efficient generative learning and online inference. In contrast to directly modeling long motion sequences, the primitive representation decomposes globally complex sequences into short and simple primitives, resulting in a more tractable data distribution for generative learning. This enables DART to learn high-quality primitive generation using as few as 10 diffusion steps. Moreover, the autoregressive and simple nature of the primitive representation makes it inherently suitable for fast online generations. Furthermore, motion primitives convey more interpretable semantics than individual frames, enhancing the learning of text-conditioned motion space.
Each motion primitive $\mathbf{P}^{i}=[\mathbf{H}^{i},\mathbf{X}^{i}]$ is a short motion clip containing $H$ frames of history motion $\mathbf{H}^{i}=[\mathbf{h}^{i,1},...,\mathbf{h}^{i,H}]$ that overlap with the previous motion primitive, and $F$ frames of future motion $\mathbf{X}^{i}=[\mathbf{x}^{i,1},...,\mathbf{x}^{i,F}]$. The history motion of the i-th motion primitive $\mathbf{H}^{i}$ consists of the last $H$ frames of the previous motion primitive $\mathbf{X}^{i-1,F-H+1:F}$. Therefore, infinitely long motions can be represented as the rollout of such overlapping motion primitives as $\mathbf{M}=[\mathbf{H}_{seed},\mathbf{X}^{1},...,\mathbf{X}^{N}]$.
We represent each motion frame using an overparameterized representation based on the SMPL-X (Pavlakos et al., 2019) parametric human body model. The overparameterized representation is a $D=276$ dimensional vector including the body root translation $\mathbf{t}$, root orientation $\mathbf{R}$, local joint rotations $\mathbf{\theta}$, joint locations $\mathbf{J}$, and the temporal difference features of locations and rotations. All the frames within one motion primitive are canonicalized in a local coordinate frame centered at the first-frame body of the primitive. We use history length $H=2$ and future length $F=8$ in our experiments. Further details of the primitive representation are attached in the Appendix A.
@@ 3.2 DART: A Diffusion-Based Autoregressive Motion Primitive Model
Figure 1: Architecture illustration of DART. The encoder network compresses the future frames $\mathbf{X}=[\mathbf{x}^{1},...,\mathbf{x}^{F}]$ into a latent variable, conditioned on the history frames $\mathbf{H}=[\mathbf{h}^{1},...,\mathbf{h}^{H}]$. The decoder network reconstructs the future frames conditioned on the history frames and latent sample $z$. The denoiser network predicts the clean latent sample $\hat{z}_{0}$ conditioned on the noising step, text prompt, history frames, and noised latent sample $z_{t}$. During the denoiser training, both the encoder and decoder networks remain frozen.
@@
We propose a novel latent diffusion model (Rombach et al., 2022; Chen et al., 2023) designed for motion primitives. This model can autoregressively generate motions with smooth transitions, conditioned on text prompts and motion history. The proposed model contains a variational autoencoder (VAE) (Kingma & Welling, 2022) that compresses the motion primitives into a compact latent space, and a latent denoising diffusion model that predicts clean latent variables from noises, conditioned on text prompts and motion history.
@@
Latent Motion Primitive Space. We introduce a motion primitive VAE that compresses motion primitives into a compact latent space, upon which we train our latent diffusion models, rather than using the raw motion space (Rombach et al., 2022; Chen et al., 2023). This design is inspired by the observations that raw motion data from motion capture datasets (Mahmood et al., 2019) often contain various levels of high-frequency noises, and training diffusion models on raw motion space leads to results inheriting such noise artifacts. This is evidenced by the significantly higher jittering in generated motions of the ablative model without VAE in Appendix. E.3. The compression achieved through the motion primitive VAE significantly mitigates the impact of high-frequency noise. The resulting latent representation is not only more compact but also more computationally efficient than the raw motion data, thereby enhancing the efficiency of our generative model and improving the control capabilities within the latent space.
@@
Our motion primitive VAE employs a transformer-based architecture based on Chen et al. (2023), comprising an encoder $\mathcal{E}$ and decoder $\mathcal{D}$, as shown in Fig. 1. The encoder takes as input the history motion frames $\mathbf{H}$ and future motion frames $\mathbf{X}$ as well as the learnable distribution tokens $T_{\mu}$ and $T_{\sigma}$, which are responsible for predicting the latent distribution. The latent sample $z$ is obtained from the predicted distribution via reparameterization (Kingma & Welling, 2022). The decoder then predicts the future frames $\mathbf{\hat{X}}$ from zero tokens conditioned on the latent sample $z$ and the history frames $\mathbf{H}$. The motion primitive VAE is trained with the future frame reconstruction loss $L_{rec}$, auxiliary losses $L_{aux}$ that penalize unnatural motion reconstruction, and a small Kullback-Leibler (KL) regularization term $L_{KL}$. We refer to Appendix C for further details about the motion primitive VAE.

### v2

# DartControl: A Diffusion-Based Autoregressive Motion Model for Real-Time Text-Driven Motion Control

@@ Abstract
Text-conditioned human motion generation, which allows for user interaction through natural language, has become increasingly popular. Existing methods typically generate short, isolated motions based on a single input sentence. However, human motions are continuous and extend over long periods, carrying rich semantics. Creating long, complex motions that precisely respond to streams of text descriptions, particularly in an online and real-time setting, remains a significant challenge. Furthermore, incorporating spatial constraints into text-conditioned motion generation presents additional challenges, as it requires aligning the motion semantics specified by text descriptions with geometric information, such as goal locations and 3D scene geometry. To address these limitations, we propose DartControl, in short DART, a Diffusion-based Autoregressive motion primitive model for Real-time Text-driven motion Control. Our model effectively learns a compact motion primitive space jointly conditioned on motion history and text inputs using latent diffusion models. By autoregressively generating motion primitives based on the preceding motion history and current text input, DART enables real-time, sequential motion generation driven by natural language descriptions. Additionally, the learned motion primitive space allows for precise spatial motion control, which we formulate either as a latent noise optimization problem or as a Markov decision process addressed through reinforcement learning. We present effective algorithms for both approaches, demonstrating our model’s versatility and superior performance in various motion synthesis tasks. Experiments show our method outperforms existing baselines in motion realism, efficiency, and controllability. Video results and code are available at https://zkf1997.github.io/DART/.
@@ 1 Introduction
Text-conditioned human motion generation has gained increasing popularity in recent years for flexible user interaction via natural languages. Existing text-conditioned motion models  (Tevet et al., 2023; Guo et al., 2024; Zhang et al., 2023a; Guo et al., 2022; Jiang et al., 2024a) primarily focus on generating standalone short motions from a single descriptive sentence. These methods fail to accurately generate long and complex motions composed of multiple action segments, where each segment is conditioned on distinct action descriptions. FlowMDM (Barquero et al., 2024) is the state-of-the-art temporal motion composition method, capable of generating complex, continuous motions by composing desired actions with precise adherence to their specified durations. However, FlowMDM is an offline method that requires prior knowledge of the entire action timeline and has a slow generation speed, making it unsuitable for online and real-time applications.
@@
In addition to text-based semantic control, generating human motion within spatial constraints and achieving specific goals, such as reaching a keyframe body pose, following a joint trajectory, or interacting with objects, has broad applications but introduces additional complex challenges. Recent works (Shafir et al., 2024; Karunratanakul et al., 2024b; Xie et al., 2024) have sought to integrate text-conditioned motion models with spatial control capabilities. However, they still face challenges in effectively balancing spatial control, motion quality, and semantic alignment with text. Moreover, these approaches are typically restricted to controlling isolated short motions in an offline setting. Meanwhile, interactive character control (Kovar et al., 2008; Holden et al., 2015; Ling et al., 2020; Peng et al., 2022) has been a longstanding focus in computer graphics, with a primary emphasis on achieving motion realism and responsiveness to interactive control signals. However, most of these works lack support for text-conditioned semantic control and are limited by being trained on small, curated datasets. Incorporating text-conditioned motion generation could provide a novel intuitive language interface for animators and everyday users to control the characters, reducing the effort required when specifying detailed spatial control signals is challenging or tedious.
@@
To address these limitations, we propose DART, a diffusion-based autoregressive motion primitive model for real-time text-driven motion control. Moreover, the compact and expressive motion space of DART provides a foundation for integrating precise spatial control through latent space optimization or reinforcement learning (RL)-based policies. DART features three key components.
@@
First, DART represents human motion as a collection of motion primitives (Zhang & Tang, 2022), which are autoregressive representations consisting of overlapping short motion segments tailored for online generation and control. These short primitives also provide a more precise alignment with atomic action semantics compared to longer sequences, enabling effective learning of a text-conditioned motion space. By focusing on shorter primitives, DART avoids the complexities and extensive data demands of modeling entire motion sequences, allowing for high-quality motion generation with only a few diffusion steps.
@@
Next, DART learns a text-conditioned autoregressive motion generation model from large-scale data using a latent diffusion architecture, which contains a variational autoencoder for learning a compact latent motion primitive space, and a denoiser network for generating motion primitives conditioned on texts and history. Leveraging the trained denoiser and decoder models, DART employs an autoregressive rollout to synthesize motion sequences from real-time text inputs, enabling the efficient generation of motions of arbitrary length. Compared with the offline temporal motion composition method FlowMDM, DART provides real-time response and 10x generation speed.
@@
Lastly, we introduce a latent space control framework based on DART for spatially controllable motion synthesis, leveraging its learned space of realistic human motions to ensure high-quality generation. We present effective optimization and learning algorithms that explore the latent diffusion noise space to synthesize motion sequences that precisely follow both textual and spatial constraints. We evaluate DART across various motion synthesis tasks, including generating long, continuous sequences from sequential text prompts, in-between motion generation, scene-conditioned motion, and goal-reaching synthesis. The experimental results show that DART is a simple, unified and highly effective motion model, consistently outperforming or matching the performance of baselines.
@@ 2 Related Works
Conditional Motion Generation. Generating realistic and diverse human motions is a long-standing challenge in computer vision and graphics. Apart from generating highly realistic human motions (Kovar et al., 2008; Holden et al., 2020; Clavet et al., 2016; Zinno, 2019), conditional generation is another important factor that aligns motion generation with human intentions and various application constraints. Text-conditioned motion generation (Tevet et al., 2023; Zhang et al., 2022a; Petrovich et al., 2022; Guo et al., 2022; Jiang et al., 2024a; Zhang et al., 2023a) has become increasingly popular since it allows users to modulate motion generation with flexible natural languages. Audio and speech-driven motion synthesis methods (Alexanderson et al., 2023; Tseng et al., 2023; Siyao et al., 2022; Ao et al., 2022; 2023) have also made significant progress recently. Moreover, there exist many applications that require spatial awareness and precise control in motion generation, such as interactive character control (Ling et al., 2020; Peng et al., 2022; Starke et al., 2022; Luo et al., 2024; Starke et al., 2024), human-scene interactions(Hassan et al., 2021; Starke et al., 2019; Zhao et al., 2022; 2023; Li et al., 2024a; Xu et al., 2023; Jiang et al., 2024b; Wang et al., 2024; Liu et al., 2024; Li et al., 2024b; Zhang et al., 2022b; 2025), and human-human(noid) interactions (Liang et al., 2024; Zhang et al., 2023c; Christen et al., 2023; Cheng et al., 2024; Shan et al., 2024). Synthesizing high-quality motions with precise spatial control remains challenging, and DART is a step toward a general and efficient motion model that supports precise control tasks.
@@
Diffusion Generative Models. Denoising Diffusion Models (Ho et al., 2020; Song et al., 2021a; b) are generative models that learn to predict clean data samples by iteratively annealing the noise from a standard Gaussian sample. Diffusion models have achieved unprecedented performances in many generation tasks including images, videos, and 3D human motions (Tevet et al., 2023; Rombach et al., 2022; Ho et al., 2022). Diffusion models can accept flexible conditions to modulate the generation, such as text prompts, images, audio, and 3D objects (Rombach et al., 2022; Tevet et al., 2023; Alexanderson et al., 2023; Tseng et al., 2023; Zhang et al., 2023b; Xu et al., 2023; Li et al., 2023). Most existing diffusion-based motion generation methods focus on offline generations of short, isolated motion sequences while neglecting the autoregressive nature of human motions (Tevet et al., 2023; Barquero et al., 2024; Chen et al., 2023b; Karunratanakul et al., 2024b; Cohan et al., 2024; Dai et al., 2025; Chen et al., 2023a). Among these methods, DNO (Karunratanakul et al., 2024b) is closely related to our optimization-based control approach, as both use diffusion noises as the latent space for editing and control. However, the key distinction lies in our latent motion primitive-based diffusion model, which, unlike DNO’s diffusion model trained with full motion sequences in explicit motion representations, achieves superior performance in harmonizing spatial control with text semantic alignment during experiments. There are also works that incorporate history conditions in diffusion-based motion generation and capture (Xu et al., 2023; Jiang et al., 2024b; Van Wouwe et al., 2024; Han et al., 2024). Shi et al. (2024) and Chen et al. (2024) adapt diffusion models for real-time character motion generation and control. While Shi et al. (2024) learns character control policies using diffusion noises as the action space, akin to our reinforcement-learning-based control, their method focuses on single-frame autoregressive generation and lacks support for text conditions, which offers a compact and intuitive interface for users to control character behaviors. In contrast, DART is an efficient and general motion model that scales effectively to large motion-text datasets. DART supports natural language interfacing and provides a versatile foundation for various motion generation tasks with spatial control. Concurrent work, CloSD (Tevet et al., 2025), trains autoregressive motion diffusion models conditioned on target joint locations to guide the human motion towards these goals. This approach relies on paired training data of control signals and human motions. In contrast, DART learns a latent motion space from motion-only data and introduces latent space control methods to achieve flexible control goals without the need for paired training data.
@@ 3.1 Preliminaries
Autoregressive Motion Primitive Representation. We model long-term human motions as the sequential composition of motion primitives (Zhang & Tang, 2022) with overlaps for efficient generative learning and online inference. Each motion primitive $\mathbf{P}^{i}=[\mathbf{H}^{i},\mathbf{X}^{i}]$ is a short motion clip containing $H$ frames of history motion $\mathbf{H}^{i}=[\mathbf{h}^{i,1},...,\mathbf{h}^{i,H}]$ that overlap with the previous motion primitive, and $F$ frames of future motion $\mathbf{X}^{i}=[\mathbf{x}^{i,1},...,\mathbf{x}^{i,F}]$. The history motion of the i-th motion primitive $\mathbf{H}^{i}$ consists of the last $H$ frames of the previous motion primitive $\mathbf{X}^{i-1,F-H+1:F}$. Therefore, infinitely long motions can be represented as the rollout of such overlapping motion primitives as $\mathbf{M}=[\mathbf{H}_{seed},\mathbf{X}^{1},...,\mathbf{X}^{N}]$. To represent human bodies in each motion frame, we use an overparameterized representation based on the SMPL-X (Pavlakos et al., 2019) parametric human body model. Each frame is represented as a $D=276$ dimensional vector including the body root translation $\mathbf{t}$, root orientation $\mathbf{R}$, local joint rotations $\bm{\theta}$, joint locations $\mathbf{J}$, and the temporal difference features of locations and rotations. Each motion primitive is canonicalized in a local coordinate frame centered at the first-frame body pelvis. We use history length $H=2$ and future length $F=8$ in our experiments. Further details of the primitive representation are attached in the Appendix A. For brevity, we omit the primitive index superscript $i$ when discussing in the context of a single primitive.
In contrast to directly modeling long motion sequences, the primitive representation decomposes globally complex sequences into short and simple primitives, resulting in a more tractable data distribution for generative learning. Moreover, the autoregressive and simple nature of the primitive representation makes it inherently suitable for fast online generations. Furthermore, motion primitives convey more interpretable semantics than individual frames, enhancing the learning of text-conditioned motion space.
@@ 3.2 DART: A Diffusion-Based Autoregressive Motion Primitive Model
Figure 1: Architecture illustration of DART. The encoder network compresses the future frames $\mathbf{X}=[\mathbf{x}^{1},...,\mathbf{x}^{F}]$ into a latent variable, conditioned on the history frames $\mathbf{H}=[\mathbf{h}^{1},...,\mathbf{h}^{H}]$. The decoder network reconstructs the future frames conditioned on the history frames and the latent sample. The denoiser network predicts the clean latent sample $\hat{\mathbf{z}}_{0}$ conditioned on the noising step, text prompt, history frames, and noised latent sample $\mathbf{z}_{t}$. During the denoiser training, the encoder and decoder network weights remain fixed.
@@
We propose a latent diffusion model (Rombach et al., 2022; Chen et al., 2023b) designed for seamless autoregressive motion generation, conditioned on text prompts and motion history. The proposed model contains a variational autoencoder (VAE) (Kingma & Welling, 2014) that compresses the motion primitives into a compact latent space and a latent denoising diffusion model that predicts clean latent variables from noise, conditioned on text prompts and motion history.
@@
Learning the Latent Motion Primitive Space. We introduce a motion primitive VAE that compresses motion primitives into a compact latent space, upon which we train our latent diffusion models, rather than using the raw motion space (Rombach et al., 2022; Chen et al., 2023b). The design of learning a compressed latent space of motion primitives is inspired by the observations that raw motion data from the used motion capture dataset AMASS (Mahmood et al., 2019) often contain various levels of artifacts including glitches and jitters, and training diffusion models on raw motion space leads to results inheriting such artifacts. This is evidenced by the significantly higher jittering in generated motions of the ablative model without VAE in Appendix. E.3. The compression achieved through the motion primitive VAE significantly mitigates the impacts of these outlier artifacts in motion data. The resulting latent representation is not only more compact but also more computationally efficient than the raw motion data, thereby enhancing the efficiency of our generative model and improving the control capabilities within the latent space.
@@
Our motion primitive VAE employs a transformer-based architecture based on MLD (Chen et al., 2023b), comprising an encoder $\mathcal{E}$ and decoder $\mathcal{D}$, as shown in Fig. 1. The encoder takes as input the history motion frames $\mathbf{H}$ and future motion frames $\mathbf{X}$ as well as the learnable distribution tokens $\mathbf{T}_{\mu}$ and $\mathbf{T}_{\sigma}$, which are responsible for predicting the distribution mean and variance. The latent sample $\mathbf{z}$ is obtained from the predicted distribution via reparameterization (Kingma & Welling, 2014). The decoder then predicts the future frames $\mathbf{\hat{X}}$ from zero tokens conditioned on the latent sample $\mathbf{z}$ and the history frames $\mathbf{H}$. The motion primitive VAE is trained with the future frame reconstruction loss $L_{rec}$, auxiliary losses $L_{aux}$ that penalize unnatural motion reconstruction, a small Kullback-Leibler (KL) regularization term $L_{KL}$, and additional SMPL-based reconstruction and regularization terms. We refer to Appendix C for further details about the motion primitive VAE.
Latent Motion Primitive Diffusion Model. Building on the compact latent motion primitive space, we formulate text-conditioned autoregressive motion generation via a probabilistic distribution $q(\mathbf{z}|\mathbf{H},c)$, and train a latent diffusion model $\mathcal{G}$ to approximate it. Diffusion models (Ho et al., 2020) are generative models that learn data distributions by progressively reversing a forward diffusion process, which iteratively adds Gaussian noise to data samples until they approach pure noise $\mathcal{N}(\mathbf{0},\mathbf{I})$. Given a motion primitive sampled from the dataset and its latent representation $\mathbf{z}_{0}$ obtained using the encoder $\mathcal{E}$, the forward diffusion produces a sequence of increasingly noisy samples $\mathbf{z}_{1},\dots,\mathbf{z}_{T}$ by iteratively adding noises as: $q(\mathbf{z}_{t}|\mathbf{z}_{t-1})=\mathcal{N}(\sqrt{1-\beta_{t}}\mathbf{z}_{t-1},\beta_{t}\mathbf{I})$, where $\beta_{t}$ are noise schedule hyper parameters. The denoiser model learns the reverse process $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},t,\mathbf{H},c)=\mathcal{N}(\bm{\mu}_{t},\bm{\Sigma}_{t})$ for generating motion primitives conditioned on the motion history $\mathbf{H}$ and text label $c$ of the motion primitive. The vairance $\bm{\Sigma}_{t}$ of the reverse process distribution is scheduled using hyper parameters, while the mean $\bm{\mu}_{t}$ is modeled using a denoiser neural network. We design the denoiser model $\mathcal{G}$ to predict the clean latent variable $\hat{\mathbf{z}}_{0}=\mathcal{G}(\mathbf{z}_{t},t,\mathbf{H},c)$, from which $\bm{\mu}_{t}$ can be derived as follows:
$\bm{\mu}_{t}=\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\mathcal{G}(\mathbf{z}_{t},t,\mathbf{H},c)+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\mathbf{z}_{t},$ (1)
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. During generation, we initialize with $\mathbf{z}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and use the denoiser model to predict the clean variable $\hat{\mathbf{z}}_{0}$, which is subsequently diffused to a lower noise level $\mathbf{z}_{T-1}$.This denoising process iterates until a clean sample $\mathbf{z}_{0}$ is obtained.